OpenStudy (anonymous):

Order 34 x 102, 1.2 x 107, 8.11 x 10-3, and 435 from least to greatest.

OpenStudy (anonymous):

34*102=3468 1.2*107=128.4 8.11*10-3=78.1 435 8.11*10-3, 1.2*107, 435, 34*102

OpenStudy (anonymous):

what is 10^-3

OpenStudy (anonymous):

ohh my bad, that will be 8110 1.2*107, 435, 34*102, 8.11*10-3